# Finite-Sample Analysis of Policy Evaluation for Robust Average Reward Reinforcement Learning

### v2

@@ C.2 Proof of Theorem 2
We follow similar process for constructing our desired semi-norm $\|\cdot\|_{\mathcal{P}}$ as in Appendix C.1 by first constructing a norm such that all $Q_{\mathsf{P}}^{\pi}$ are strictly less then one under that norm. We choose $\alpha\in(r^{*},1)$ and we follow the approach in Wirth (2002) by constructing an extremal norm $\|\cdot\|_{\rm ext}$ as follows:
@@
Lemma 6.
The operator $\|\cdot\|_{\rm ext}$ is a valid norm with $\|Q_{\mathsf{P}}^{\pi}\|_{\rm ext}<1$ for all $\mathsf{P}\in\mathcal{P}$
@@
We first prove that $\|\cdot\|_{\rm ext}$ is bounded. Following Lemma 10 and choosing $\lambda\in(r^{*},\alpha)$, then there exist a positive constant $C<\infty$ such that
@@
To check that $\|\cdot\|_{\rm ext}$ is a valid norm, note that if $x=\bf{0}$, $\|x\|_{\rm ext}$ is directly 00. On the other hand, if $\|x\|_{\rm ext}=0$, we have
@@
Since $\mathsf{P}$ is arbitrary, (6) implies that for any $\mathsf{P}\in\mathcal{P}$,
@@
Lemma 7.
The operator $\|\cdot\|_{\mathcal{P}}$ is a valid semi-norm with kernel being exactly $\{c\mathbf{e}:c\in\mathbb{R}\}$. Furthermore, for all $x\in\mathbb{R}^{S}$, we have $\bigl{\|}\mathsf{P}^{\pi}x\bigr{\|}_{\mathcal{P}}\leq(\alpha+\epsilon)\bigl{\|}x\bigr{\|}_{\mathcal{P}}$.
@@
Regarding the kernel, if $x=k\mathbf{e}$ for some $k\in\mathbb{R}$, then similar to (5), we have
@@
Since $\alpha\in(0,1)$ and $\epsilon\in(0,1-\alpha)$. Thus, let $\gamma=\alpha+\epsilon$ then $\gamma\in(0,1)$. Substituting the above result back to (C.2), we obtain
@@
Appendix D Biased Stochastic Approximation Convergence Rate
In Section 3, we established that the robust Bellman operator is a contraction under the semi-norm $\|\cdot\|_{\mathcal{P}}$, ensuring that policy evaluation can be analyzed within a well-posed stochastic approximation framework. However, conventional stochastic approximation methods typically assume unbiased noise, where variance diminishes over time without introducing systematic drift. In contrast, the noise in robust policy evaluation under TV and Wasserstein distance uncertainty sets exhibits a small but persistent bias, arising from the estimators of the support functions $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$ (discussed in Section 4). This bias, if not properly addressed, can lead to uncontrolled error accumulation, affecting the reliability of policy evaluation. To address this challenge, this section introduces a novel analysis of biased stochastic approximation, leveraging properties of dual norms to ensure that the bias remains controlled and does not significantly impact the convergence rate. Our results extend prior work on unbiased settings and provide the first explicit finite-time guarantees, which is further used to establish the sample complexity of policy evaluation in robust average reward RL. Specifically, we analyze the iteration complexity for solving the fixed equivalent class equation $H(x^{*})-x^{*}\in\overline{E}$ where $\overline{E}\coloneqq\{c\mathbf{e}:c\in\mathbb{R}\}$ with $\mathbf{e}$ being the all-ones vector. The stochastic approximation iteration being used is as follows:
@@ Appendix D Biased Stochastic Approximation Convergence Rate
Theorem 8 adapts the analysis of (Zhang et al., 2021) and extends it to a biased i.i.d. noise setting. To manage the bias terms, we leverage properties of dual norms to bound the inner product between the error term and the gradient, ensuring that the bias influence remains logarithmic in $T$ rather than growing unbounded, while also carefully structuring the stepsize decay to mitigate long-term accumulation. This results in an extra $\varepsilon_{\mathrm{bias}}$ term with logarithmic dependence of the total iteration $T$.
We perform analysis of the biased‐noise extension to the semi-norm stochastic approximation (SA) problem by constructing a smooth convex semi-Lyapunov function for forming the negative drift Zhang et al. (2021); Chen et al. (2025) and using properties in dual norms for managing the bias.
D.1 Proof of Theorem 8
D.1.1 Setup and Notation.
In this section, we override the notation of the semi-norm $\|\cdot\|_{\mathcal{P}}$ by re-writing it as the norm $\|\cdot\|_{\mathcal{N}}$ (defined in (64)) to the equivalence class of constant vectors. For any norm $\|\cdot\|_{c}$ and equivalent class $\delta_{\overline{E}}$, define the indicator function $\delta_{\overline{E}}$ as
@@ D.1 Proof of Theorem 8
then by Zhang et al. (2021), the semi-norm induced by norm $\|\cdot\|_{c}$ and equivalent class $\overline{E}$ is the infimal convolution of $\|\cdot\|_{c}$ and the indicator function $\delta_{\overline{E}}$ can be defined as follows
$\|x\|_{c,\overline{E}}\coloneqq(\|\cdot\|_{c}\ast_{\inf}\delta_{\overline{E}})(x)=\inf_{y}(\|x-y\|_{c}+\delta_{\overline{E}}(y))=\inf_{e\in\overline{E}}\|x-e\|_{c}\quad\forall x.$ (62)
@@
$\|x\|_{\mathcal{N}}\geq\sup_{\mathsf{P}\in\mathcal{P}}\bigl{\|}Q_{\mathsf{P}}^{\pi}x\bigr{\|}_{\rm ext}+\epsilon\|x\|_{\rm ext}>0.$
@@
We thus restate our problem of analyzing the iteration complexity for solving the fixed equivalent class equation $H(x^{*})-x^{*}\in\overline{E}$, with the operator $H:\mathbb{R}^{n}\to\mathbb{R}^{n}$ satisfying the contraction property as follows:
@@
  • $\mathbb{E}[\,\|w^{t}\|_{\mathcal{N},\overline{E}}^{2}|\mathcal{F}^{t}]\leq A+B\|x^{t}-x^{*}\|_{\mathcal{N},\overline{E}}^{2}$ (In the robust average reward TD case, $B=0$).
@@
  • $\eta_{t}>0$ is a chosen step‐size sequence (decreasing or constant).
Note that beside the bias in the noise, the above formulation and assumptions are identical to the unbiased setups in Section B of Zhang et al. (2021). Thus, we emphasize mostly on managing the bias.
D.1.2 Semi‐Lyapunov $M_{\overline{E}}(\cdot)$ and Smoothness.
By (Zhang et al., 2021, Proposition 1–2), using the Moreau envelope function $M(x)$ in Definition 2.2 of Chen et al. (2020), we define
@@
D.1.3 Formal Statement of Theorem 8
By $L$‐smoothness w.r.t. $\|\cdot\|_{s,\overline{E}}$ in (68), for each $t$,
@@ D.1.3 Formal Statement of Theorem 8
[…]
@@
Combining (82) with (D.1.3) we obtain
@@
To bound the first term in the RHS of (D.1.3), note that
@@
where $(a)$ follows from the convexity of $M_{\bar{E}}$, $(b)$ follows from $x^{*}$ belonging to a fixed equivalent class with respect to $H$ and $(c)$ follows from the contraction property of $H$. Combining (D.1.3). (D.1.3) and Lemma 11 with (72), we arrive as follows:
@@
Where $\alpha_{2}\coloneqq(1-\gamma\sqrt{c_{{u}}/c_{{l}}})$, $\alpha_{3}\coloneqq(8+2B)c_{{u}}\rho_{2}L$ and $\alpha_{4}\coloneqq A\rho_{2}L$. We now present the formal version of Theorem 8 as follows:
Theorem 9 (Formal version of Theorem 8).
let $\alpha_{2}$, $\alpha_{3}$ and $\alpha_{4}$ be defined in (D.1.3), if $x^{t}$ is generated by (66) with all assumptions in D.1.1 satisfied, then if the stepsize $\eta_{t}\coloneqq\frac{1}{\alpha_{2}(t+K)}$ while $K\coloneqq\max\{\alpha_{3}/\alpha_{2},3\}$,
@@
where $C_{1}=G(1+2x_{sup})$, $C_{2}=\frac{1}{K}+\log\big{(}\frac{T-1+K}{K}\big{)}$, $G$ is defined in (81) and $x_{sup}\coloneqq\sup\|x\|_{\mathcal{N},\overline{E}}$ is the upper bound of the $\|\cdot\|_{\mathcal{P}}$ norm for all $x^{t}$.
@@
This choice $\eta_{t}$ satisfies $\alpha_{3}\eta_{t}^{2}\leq\alpha_{2}\eta_{t}$. Thus, by (D.1.3) we have
@@
where the term $R_{1}$ is identical to the unbiased case in Theorem 3 of Zhang et al. (2021) which leads to
@@
Appendix E Uncertainty Set Support Function Estimators
E.1 Proof of Theorem 3
@@
E.2 Proof of Theorem 4
denote $\hat{\sigma}^{*}_{\mathcal{P}^{a}_{s}}(V)$ as the untruncated MLMC estimator obtained by running Algorithm 1 when setting $N_{\max}$ to infinity. From Wang et al. (2023b), under both TV uncertainty sets and Wasserstein uncertainty sets, we have $\hat{\sigma}^{*}_{\mathcal{P}^{a}_{s}}(V)$ as an unbiased estimator of ${\sigma}_{\mathcal{P}^{a}_{s}}(V)$. Thus,
@@ E.2 Proof of Theorem 4
By the binomial concentration and the Lipschitz property of the support function as in Lemma 2, we know for TV distance uncertainty, we have
@@ E.3 Proof of Lemma 2
Combining (107) with (E.3) we obtain the first part of (23).
@@ E.4 Proof of Theorem 5
Appendix F Convergence for Robust TD
F.1 Formal Statement of Theorem 6
@@
where $p_{\pi}$ is the finite state Markov chain induced by $\pi$, $\mu_{0}$ is any initial probability distribution on $\mathcal{S}$ and $\nu$ is its invariant distribution. By Assumption 1, and Lemma 13,and for any value function $V$, we have
@@
where we define $t_{\mathrm{mix}}\coloneqq\sup_{p\in\mathcal{P}}t^{p}_{\mathrm{mix}}$, then $t_{\mathrm{mix}}$ is also finite due to the compactness of $\mathcal{P}$. We now derive the bounds of biases and variances for the three types of uncertainty sets. Regarding contamination uncertainty sets, according to Lemma 12, $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$ is unbiased and has variance bounded by $\|V\|^{2}$. Thu, define $t_{\mathrm{mix}}$ according to Lemma 13 and combining the above result with Lemma 13, we obtain that $\hat{\mathbf{T}}_{g_{0}}(V_{t})$ is also unbiased and the variance satisfies
@@
Regarding TV distance uncertainty sets, using the property of the bias and variance of $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$ in Theorem 4 and Theorem 5 while combining them with Lemma 13, we have
@@
Since $x_{i}=0$, for every $j$ we have $-\|x\|_{\infty}\leq x_{j}\leq\|x\|_{\infty}$. Hence
@@
With the relationship established above, line 1 - line 7 of Algorithm 2 can be formally treated as a special instance of the SA updates in (66) with $B=0$. We now provide the bias and variance of the i.i.d. noise for the different uncertainty sets discussed using Lemma 9 and the estimation bounds in (121)-(F.1). For contamination uncertainty sets, we have
@@
Theorem 10 (Formal version of Theorem 6).
Let $\alpha_{2}\coloneqq(1-\gamma\sqrt{c_{{u}}/c_{{l}}})$, $\alpha_{3}\coloneqq 8c_{{u}}\rho_{2}L$ and $\alpha_{4}\coloneqq\rho_{2}L$, if $V_{t}$ is generated by Algorithm 2. Define $V^{*}$ to be the anchored robust value function $V^{*}=V^{\pi}_{\mathsf{P}_{V}}+c\mathbf{e}$ for some $c$ such that $V^{*}(s_{0})=0$, then under Assumption 1 and if the stepsize $\eta_{t}\coloneqq\frac{1}{\alpha_{2}(t+K)}$ while $K\coloneqq\max\{\alpha_{3}/\alpha_{2},3\}$, then for contamination uncertainty sets,
@@
where the $\varepsilon$ and $A$ terms are defined in (127)-(F.1), $C_{2}=\frac{1}{K}+\log\big{(}\frac{T-1+K}{K}\big{)}$, $C_{3}=G(1+8C_{\mathcal{P}}t_{\mathrm{mix}})$, $\gamma$ is defined in (14), $c_{u},c_{l}$ are defined in (67), $\rho_{2}$ is defined in (70), $G$ is defined in (81), and $C_{\mathcal{P}},c_{\mathcal{P}}$ are defined in Lemma 9.
@@
By Lemma 9 and (120), we have that for any value function $V$ its $\|\cdot\|_{\mathcal{P}}$ norm is bounded as follows:
@@
Substituting the terms of (127)-(F.1), (137), and Theorem 2 to Theorem 9, we would have for contamination uncertainty sets,
@@
where the $\varepsilon$ and $A$ terms are defined in (127)-(F.1), $C_{2}=\frac{1}{K}+\log\big{(}\frac{T-1+K}{K}\big{)}$, $C_{3}=G(1+8C_{\mathcal{P}}t_{\mathrm{mix}})$, $\gamma$ is defined in (14), $c_{u},c_{l}$ are defined in (67), $\rho_{2}$ is defined in (70), $G$ is defined in (81), and $C_{\mathcal{P}}$ is defined in Lemma 9. We now translate the result back to the standard $l_{\infty}$ norm by applying Lemma 9 again to thee above, we obtain the desired results. ∎
F.2 Proof of Theorem 6
We use the result from Theorem 10, to set $\mathbb{E}\Bigl{[}\|V_{T}-V^{*}\|^{2}_{\infty}\Bigr{]}\leq\epsilon^{2}$. For contamination uncertainty set we set $T=\operatorname*{\mathcal{O}}\left(\frac{t^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\right)$, resulting in $\operatorname*{\mathcal{O}}\left(\frac{SAt^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\right)$ sample complexity. For TV and Wasserstein uncertainty set, we set $N_{\max}=\operatorname*{\mathcal{O}}\left(\log\frac{t_{\mathrm{mix}}}{\epsilon(1-\gamma)}\right)$ and $T=\operatorname*{\mathcal{O}}\left(\frac{t^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\log\frac{t_{\mathrm{mix}}}{\epsilon(1-\gamma)}\right)$, combining with Theorem 3, this would result in $\operatorname*{\mathcal{O}}\left(\frac{SAt^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\log^{2}\frac{t_{\mathrm{mix}}}{\epsilon(1-\gamma)}\right)$ sample complexity.
[…]
F.3 Formal Statement of Theorem 7
@@
By the Bellman equation in Theorem 1, we have $g^{\pi}_{\mathcal{P}}=\bar{\delta}(V^{*})$, which implies
@@
Where the last inequality is by Lemma 9. Thus, the following recursion can be formed

### v3

@@ A.2.3 Extremal Norm Construction
We follow similar process for constructing our desired semi-norm $\|\cdot\|_{\mathcal{P}}$ as in Appendix A.1 by first constructing a norm such that all $Q_{\mathsf{P}}^{\pi}$ are strictly less then one under that norm. We choose $\alpha\in(r^{*},1)$ and we follow the approach in [40] by constructing an extremal norm $\|\cdot\|_{\rm ext}$ as follows:
@@
Lemma A.7.
Under Assumption 3.1 and the radius conditions of Lemma A.4-A.6, the operator $\|\cdot\|_{\rm ext}$ is a valid norm with $\|Q_{\mathsf{P}}^{\pi}\|_{\rm ext}<1$ for all $\mathsf{P}\in\mathcal{P}$
@@
We first prove that $\|\cdot\|_{\rm ext}$ is bounded. Following Lemma F.1 and choosing $\lambda\in(r^{*},\alpha)$, then there exist a positive constant $C<\infty$ such that
@@
To check that $\|\cdot\|_{\rm ext}$ is a valid norm, note that if $x=\bf{0}$, $\|x\|_{\rm ext}$ is directly 0. On the other hand, if $\|x\|_{\rm ext}=0$, we have
@@
Since $\mathsf{P}$ is arbitrary, (A.7) implies that for any $\mathsf{P}\in\mathcal{P}$,
@@
A.2.4 Semi-Norm Contraction for Robust Bellman Operator
@@
Lemma A.8.
The operator $\|\cdot\|_{\mathcal{P}}$ is a valid semi-norm with kernel being exactly $\{c\mathbf{e}:c\in\mathbb{R}\}$ under Assumption 3.1 and the radius conditions of Lemma A.4-A.6. Furthermore, for all $x\in\mathbb{R}^{S}$, we have $\bigl\|\mathsf{P}^{\pi}x\bigr\|_{\mathcal{P}}\leq(\alpha+\epsilon)\bigl\|x\bigr\|_{\mathcal{P}}$ for all $\mathsf{P}\in\mathcal{P}$.
@@
Regarding the kernel, if $x=k\mathbf{e}$ for some $k\in\mathbb{R}$, then similar to (A.2), we have
@@
Since $\alpha\in(0,1)$ and $\epsilon\in(0,1-\alpha)$. Thus, let $\gamma=\alpha+\epsilon$ then $\gamma\in(0,1)$. Substituting the above result back to (A.2), we obtain
@@
Appendix B Biased Stochastic Approximation Convergence Rate
In Section 4, we established that the robust Bellman operator is a contraction under the semi-norm $\|\cdot\|_{\mathcal{P}}$, ensuring that policy evaluation can be analyzed within a well-posed stochastic approximation framework. However, conventional stochastic approximation methods typically assume unbiased noise, where variance diminishes over time without introducing systematic drift. In contrast, the noise in robust policy evaluation under TV and Wasserstein distance uncertainty sets exhibits a small but persistent bias, arising from the estimators of the support functions $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$ (discussed in Section 5). This bias, if not properly addressed, can lead to uncontrolled error accumulation, affecting the reliability of policy evaluation. To address this challenge, this section introduces a novel analysis of biased stochastic approximation, leveraging properties of dual norms to ensure that the bias remains controlled and does not significantly impact the convergence rate. Our results extend prior work on unbiased settings and provide the first explicit finite-time guarantees, which are further used to establish the sample complexity of policy evaluation in robust average-reward RL. Specifically, we analyze the iteration complexity for solving the fixed equivalence class equation $H(x^{*})-x^{*}\in\overline{E}$ where $\overline{E}\coloneqq\{c\mathbf{e}:c\in\mathbb{R}\}$ with $\mathbf{e}$ being the all-ones vector. The stochastic approximation iteration being used is as follows:
@@ Appendix B Biased Stochastic Approximation Convergence Rate
Theorem B.1 adapts the analysis of [42] and extends it to a biased i.i.d. noise setting. To manage the bias terms, we leverage properties of dual norms to bound the inner product between the error term and the gradient, ensuring that the bias influence remains logarithmic in $T$ rather than growing unbounded, while also carefully structuring the stepsize decay to mitigate long-term accumulation. This results in an extra $\varepsilon_{\mathrm{bias}}$ term with logarithmic dependence of the total iteration $T$.
We perform analysis of the biased‐noise extension to the semi-norm stochastic approximation (SA) problem by constructing a smooth convex semi-Lyapunov function for forming the negative drift [42, 8] and using properties in dual norms for managing the bias.
B.1 Proof of Theorem B.1
B.1.1 Setup and Notation.
In this section, we override the notation of the semi-norm $\|\cdot\|_{\mathcal{P}}$ by re-writing it as the norm $\|\cdot\|_{\mathcal{N}}$ (defined in (96)) to the equivalence class of constant vectors. For any norm $\|\cdot\|_{c}$ and equivalence class $\overline{E}$}, define the indicator function $\delta_{\overline{E}}$ as
@@
then by [42], the semi-norm induced by norm $\|\cdot\|_{c}$ and equivalence class $\overline{E}$ is the infimal convolution of $\|\cdot\|_{c}$ and the indicator function $\delta_{\overline{E}}$ can be defined as follows
$\|x\|_{c,\overline{E}}\coloneqq(\|\cdot\|_{c}\ast_{\inf}\delta_{\overline{E}})(x)=\inf_{y}(\|x-y\|_{c}+\delta_{\overline{E}}(y))=\inf_{e\in\overline{E}}\|x-e\|_{c}\quad\forall x,$ (94)
@@
$\|x\|_{\mathcal{N}}\geq\epsilon\|x\|_{\rm ext}>0.$
@@
We thus restate our problem of analyzing the iteration complexity for solving the fixed equivalence class equation $H(x^{*})-x^{*}\in\overline{E}$, with the operator $H:\mathbb{R}^{n}\to\mathbb{R}^{n}$ satisfying the contraction property as follows:
@@
  • $\mathbb{E}[\,\|w^{t}\|_{\mathcal{N},\overline{E}}^{2}|\mathcal{F}^{t}]\leq A+B\|x^{t}-x^{*}\|_{\mathcal{N},\overline{E}}^{2}$ (In the robust average-reward TD case, $B=0$).
@@
  • $\eta_{t}>0$ is a chosen stepsize sequence (decreasing or constant).
Note that beside the bias in the noise, the above formulation and assumptions are identical to the unbiased setups in Section B of [42]. Thus, we emphasize mostly on managing the bias.
B.1.2 Semi‐Lyapunov $M_{\overline{E}}(\cdot)$ and Smoothness.
By [42, Proposition 1–2], using the Moreau envelope function $M(x)$ in Definition 2.2 of [7], we define
@@
B.1.3 Formal Statement of Theorem B.1
By LL‐smoothness w.r.t. $\|\cdot\|_{s,\overline{E}}$ in (100), for each $t$,
@@
To analyze the additional bias term $\langle\nabla M_{\overline{E}}(x^{t}-x^{*}),\mathbb{E}[\omega^{t}|\mathcal{F}^{t}]\rangle$, we use the fact that for any (semi-)norm $\|\cdot\|$ with dual (semi-)norm $\|\cdot\|_{*}$ (defined by $\|u\|_{*}=\sup\{\langle u,v\rangle:\|v\|\leq 1\}$), we have the general inequality
@@
Combining (114) with (B.1.3) we obtain
@@
To bound the first term in the RHS of (B.1.3), note that
@@
where $(a)$ follows from the convexity of $M_{\bar{E}}$, $(b)$ follows from $x^{*}$ belonging to a fixed equivalence class with respect to $H$ and $(c)$ follows from the contraction property of $H$. Combining (B.1.3). (B.1.3) and Lemma F.2 with (104), we arrive as follows:
@@
Where $\alpha_{2}\coloneqq(1-\gamma\sqrt{c_{{u}}/c_{{l}}})$, $\alpha_{3}\coloneqq(8+2B)c_{{u}}\rho_{2}L$ and $\alpha_{4}\coloneqq A\rho_{2}L$. We now present the formal version of Theorem B.1 as follows:
Theorem B.3 (Formal version of Theorem B.1).
let $\alpha_{2}$, $\alpha_{3}$ and $\alpha_{4}$ be defined in (B.1.3), if $x^{t}$ is generated by (98) with all assumptions in B.1.1 satisfied, then if the stepsize $\eta_{t}\coloneqq\frac{1}{\alpha_{2}(t+K)}$ while $K\coloneqq\max\{\alpha_{3}/\alpha_{2},3\}$,
@@
where $C_{1}=G(1+2x_{sup})$, $C_{2}=\frac{1}{K}+\log\big(\frac{T-1+K}{K}\big)$, $G$ is defined in (113) and $x_{sup}\coloneqq\sup\|x\|_{\mathcal{N},\overline{E}}$ is the upper bound of the $\|\cdot\|_{\mathcal{P}}$ semi-norm for all $x^{t}$.
@@
This choice $\eta_{t}$ satisfies $\alpha_{3}\eta_{t}^{2}\leq\alpha_{2}\eta_{t}$. Thus, by (B.1.3) we have
@@
where the term $R_{1}$ is identical to the unbiased case in Theorem 3 of [42] which leads to
@@
Appendix C Uncertainty Set Support Function Estimators
C.1 Proof of Theorem 5.1
@@ Appendix C Uncertainty Set Support Function Estimators
C.2 Proof of Theorem 5.2
denote $\hat{\sigma}^{*}_{\mathcal{P}^{a}_{s}}(V)$ as the untruncated MLMC estimator obtained by running Algorithm 1 when setting $N_{\max}$ to infinity. From [37], under both TV uncertainty sets and Wasserstein uncertainty sets, we have $\hat{\sigma}^{*}_{\mathcal{P}^{a}_{s}}(V)$ as an unbiased estimator of ${\sigma}_{\mathcal{P}^{a}_{s}}(V)$. Thus,
@@
By the binomial concentration and the Lipschitz property of the support function as in Lemma 5.3, we know for TV distance uncertainty, we have
@@
Combining (139) with (C.3) we obtain the first part of (23).
@@
Appendix D Convergence for Robust TD
D.1 Formal Statement of Theorem 6.1
@@ Appendix D Convergence for Robust TD
where $p_{\pi}$ is the finite state Markov chain induced by $\pi$, $\mu_{0}$ is any initial probability distribution on $\mathcal{S}$ and $\nu$ is its invariant distribution. By Assumption 3.1, and Lemma F.4,and for any value function $V$, we have
@@
where we define $t_{\mathrm{mix}}\coloneqq\sup_{p\in\mathcal{P}}t^{p}_{\mathrm{mix}}$, then $t_{\mathrm{mix}}$ is also finite due to the compactness of $\mathcal{P}$. We now derive the bounds of biases and variances for the three types of uncertainty sets. Regarding contamination uncertainty sets, according to Lemma F.3, $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$ is unbiased and has variance bounded by $\|V\|^{2}$. Thu, define $t_{\mathrm{mix}}$ according to Lemma F.4 and combining the above result with Lemma F.4, we obtain that $\hat{\mathbf{T}}_{g_{0}}(V_{t})$ is also unbiased and the variance satisfies
@@
Regarding TV distance uncertainty sets, using the property of the bias and variance of $\hat{\sigma}_{\mathcal{P}^{a}_{s}}(V)$ in Theorem 5.2 and Theorem 5.4 while combining them with Lemma F.4, we have
@@
Since $x_{i}=0$, for every $j$ we have $\|x\|_{\infty}\leq x_{j}\leq\|x\|_{\infty}$. Hence
@@
With the relationship established above, line 1 - line 7 of Algorithm 2 can be formally treated as a special instance of the SA updates in (98) with $B=0$. We now provide the bias and variance of the i.i.d. noise for the different uncertainty sets discussed using Lemma D.1 and the estimation bounds in (153)-(D.1). For contamination uncertainty sets, we have
@@
Theorem D.2 (Formal version of Theorem 6.1).
Let $\alpha_{2}\coloneqq(1-\gamma\sqrt{c_{{u}}/c_{{l}}})$, $\alpha_{3}\coloneqq 8c_{{u}}\rho_{2}L$ and $\alpha_{4}\coloneqq\rho_{2}L$, if $V_{t}$ is generated by Algorithm 2. Define $V^{*}$ to be the anchored robust value function $V^{*}=V^{\pi}_{\mathsf{P}_{V}}+c\mathbf{e}$ for some $c$ such that $V^{*}(s_{0})=0$, then under Assumption 3.1 and the radius conditions of Lemma A.4-A.6, if the stepsize $\eta_{t}\coloneqq\frac{1}{\alpha_{2}(t+K)}$ while $K\coloneqq\max\{\alpha_{3}/\alpha_{2},3\}$, then for contamination uncertainty sets,
@@
where the $\varepsilon$ and $A$ terms are defined in (159)-(D.1), $C_{2}=\frac{1}{K}+\log\big(\frac{T-1+K}{K}\big)$, $C_{3}=G(1+8C_{\mathcal{P}}t_{\mathrm{mix}})$, $\gamma$ is defined in (14), $c_{u},c_{l}$ are defined in (99), $\rho_{2}$ is defined in (102), $G$ is defined in (113), and $C_{\mathcal{P}},c_{\mathcal{P}}$ are defined in Lemma D.1.
@@
By Lemma D.1 and (152), we have that for any value function $V$ its $\|\cdot\|_{\mathcal{P}}$ norm is bounded as follows:
@@
Substituting the terms of (159)-(D.1), (169), and Theorem 4.2 to Theorem B.3, we would have for contamination uncertainty sets,
@@
where the $\varepsilon$ and $A$ terms are defined in (159)-(D.1), $C_{2}=\frac{1}{K}+\log\big(\frac{T-1+K}{K}\big)$, $C_{3}=G(1+8C_{\mathcal{P}}t_{\mathrm{mix}})$, $\gamma$ is defined in (14), $c_{u},c_{l}$ are defined in (99), $\rho_{2}$ is defined in (102), $G$ is defined in (113), and $C_{\mathcal{P}}$ is defined in Lemma D.1. We now translate the result back to the standard $l_{\infty}$ norm by applying Lemma D.1 again to the above, we obtain the desired results. ∎
D.2 Proof of Theorem 6.1
We use the result from Theorem D.2, to set $\mathbb{E}\Bigl[\|V_{T}-V^{*}\|^{2}_{\infty}\Bigr]\leq\epsilon^{2}$. For contamination uncertainty set we set $T=\operatorname*{\mathcal{O}}\left(\frac{t^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\right)$, resulting in $\operatorname*{\mathcal{O}}\left(\frac{SAt^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\right)$ sample complexity. For TV and Wasserstein uncertainty set, we set $N_{\max}=\operatorname*{\mathcal{O}}\left(\log\frac{t_{\mathrm{mix}}}{\epsilon(1-\gamma)}\right)$ and $T=\operatorname*{\mathcal{O}}\left(\frac{t^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\log\frac{t_{\mathrm{mix}}}{\epsilon(1-\gamma)}\right)$, combining with Theorem 5.1, this would result in $\operatorname*{\mathcal{O}}\left(\frac{SAt^{2}_{\mathrm{mix}}}{\epsilon^{2}(1-\gamma)^{2}}\log^{2}\frac{t_{\mathrm{mix}}}{\epsilon(1-\gamma)}\right)$ sample complexity.
To show order-optimality, we provide the standard mean estimation as a hard example. Consider the TD learning of the MDP with only two states $\mathcal{S}=\{s_{1},s_{2}\}$, and $Pr(s\to s_{1})=p$, $Pr(s\to s_{2})=1-p$ for each $s\in S$ with $p\in(0,1)$. Thus, this MDP is indifferent from the actions chosen and we further define $r(s_{1})=1,\;r(s_{2})=0$. Thus, estimating the relative value functions is equivalent of estimating $p$. By the Cramér–Rao or direct variance argument for Bernoulli($p$) estimation, we have that to achieve $|\hat{p_{N}}-p|^{2}\leq\epsilon^{2}$ requires $N\geq\ 1/{2\epsilon^{2}}=\Omega(\epsilon^{-2})$.
D.3 Formal Statement of Theorem 6.2
@@
By the Bellman equation in Theorem 3.2, we have $g^{\pi}_{\mathcal{P}}=\bar{\delta}(V^{*})$, which implies
@@
Where the last inequality is by Lemma D.1. Thus, the following recursion can be formed